At the start of the current week there are 30 units of X and 90 units of Y in stock. Available processing time on machine A is forecast to be 40 hours and on machine B is forecast to be 35 hours. The demand for X in the current week is forecast to be 75 units and for Y is forecast to be 95 units. Company policy is to maximize the combined sum of the units of X and the units of Y in stock at the end of the week.

 Formulate the problem of deciding how much of each product to make in the current week as a linear program.

 Solve this linear program graphically.


1
Expert's answer
2021-11-24T14:27:55-0500

A company makes two products (X and Y) using two machines (A and B). Each unit of X that is produced requires 50 minutes processing time on machine A and 30 minutes processing time on machine B. Each unit of Y that is produced requires 24 minutes processing time on machine A and 33 minutes processing time on machine B.


Solution.

x is the number of units of X produced in the current week

y is the number of units of Y produced in the current week


constraints are:

"50x + 24y \le 40\cdot60" machine A time

"30x + 33y \le 35\cdot60" machine B time

"x \ge 75 - 30"

i.e. "x \ge 45" so production of X"\ge" demand (75) - initial stock (30), which ensures we meet demand

"y\ge 95 - 90"

i.e. "y\ge 5" so production of Y "\ge" demand (95) - initial stock (90), which ensures we meet demand


The objective is: maximise "(x+30-75) + (y+90-95) = (x+y-50)"

i.e. to maximise the number of units left in stock at the end of the week

It is plain from the diagram below that the maximum occurs at the intersection of

"x=45" and "50x + 24y = 2400"



we have that x=45 and y=6.25 with the value of the objective function being 1.25